# VLM2Vec: Training Vision-Language Models for Massive Multimodal Embedding Tasks

### v2

@@ Abstract
Embedding models have been crucial in enabling various downstream tasks such as semantic similarity, information retrieval, and clustering. Recently, there has been a surge of interest in developing universal text embedding models that can generalize across tasks (e.g., MTEB). However, progress in learning universal multimodal embedding models has been relatively slow despite their importance. In this work, we aim to explore the potential for building universal embeddings capable of handling a wide range of downstream tasks. Our contributions are twofold: (1) MMEB (Massive Multimodal Embedding Benchmark), which covers 4 meta-tasks (i.e. classification, visual question answering, multimodal retrieval, and visual grounding) and 36 datasets, including 20 training and 16 evaluation datasets, and (2) Vlm2Vec (Vision-Language Model $\rightarrow$ Vector), a contrastive training framework that converts any state-of-the-art vision-language model into an embedding model via training on MMEB. Unlike previous models such as CLIP and BLIP, Vlm2Vec can process any combination of images and text to generate a fixed-dimensional vector based on task instructions. We build a series of Vlm2Vec models on Phi-3.5-V and evaluate them on MMEB’s evaluation split. Our results show that Vlm2Vec achieves an absolute average improvement of 10% to 20% over existing multimodal embedding models on both in-distribution and out-of-distribution datasets in MMEB.
@@ 1 Introduction
- Vlm2Vec: We adopt the pre-trained vision-language model Phi-3.5-V (Abdin et al., 2024) as the backbone for Vlm2Vec. In contrast to other multimodal embedding models like UniIR (Wei et al., 2023) and MagicLens (Zhang et al., 2024), which rely on late fusion of CLIP (Radford et al., 2021) features, our approach leverages the deep integration of vision and language features within a transformer architecture. There are several advantages to this approach: (1) VLMs are trained on massive multimodal datasets and can handle any combination of images and text, as well as high-resolution images and long text inputs; (2) vision and language features are deeply fused in the transformer model, improving the model’s ability to capture cross-modal relationships; and (3) these models are well-suited for generalizing across diverse tasks, particularly those requiring instruction-following capabilities. These factors make Vlm2Vec an ideal choice for task generalization. We trained Vlm2Vec on the 20 MMEB training datasets using contrastive learning and compared its performance with various baselines.
Following extensive contrastive training, Vlm2Vec can handle any combination of images and text, producing fixed-dimensional vectors. We evaluate Vlm2Vec against a wide array of multimodal embedding models, including CLIP (Radford et al., 2021), BLIP2 (Li et al., 2023a), SigLIP (Zhai et al., 2023), MagicLens (Zhang et al., 2024), UniIR (Wei et al., 2023) and E5-V (Jiang et al., 2024), demonstrating consistent improvements across all task categories. Notably, Vlm2Vec achieves a 17.3-point improvement (from 42.8 to 60.1) across all 36 MMEB datasets and a 11.6-point increase (from 40.4 to 52.0) on 16 out-of-distribution datasets in zero-shot evaluation, highlighting the effectiveness of our proposed framework. We are committed to releasing all data, training code, and models to facilitate future research in this area.
@@ 2.1 Dataset Overview
The embedding models are supposed to compress the query side into a vector and the target candidates into a set of vectors. The candidate with the highest dot-product will be selected as the prediction for evaluation. We measure the Precision@1 to reflect the percentage of top candidate matching the groundtruth. To ensure the task difficulty, we introduce a large amount of candidates.
MMEB offers a wide range of tasks from various domains, such as common, news, Wikipedia, web, and fashion. The benchmark incorporates diverse combinations of modalities for both queries and targets, including text, images, and text-image pairs. Additionally, tasks are designed to follow different types of instructions. For instance, tasks may involve object recognition (e.g., “Identify the object shown in the image.”), retrieval (e.g., “Find an image that matches the given caption.”), or visual grounding (e.g., “Select the portion of the image that answers the question.”). Examples for each dataset in MMEB are provided in Tables 6, 7, 8 and 9. The diversity in MMEB makes it an ideal testbed for universal embeddings.
@@ 2.2 Meta-task and Dataset Design
Classification This category comprises 5 in-distribution and 5 out-of-distribution datasets. Queries consist of instructions and images, optionally accompanied by related text. Targets are class labels, and the number of class labels corresponds to the number of classes in the dataset.
Visual Question Answering This category includes 6 in-distribution and 4 out-of-distribution datasets. The query consists of an instruction, an image, and a piece of text as the question, while the target is the answer. Each query has 1,000 target candidates: 1 ground truth and 999 distractors.
Information Retrieval This category contains 8 in-distribution and 4 out-of-distribution datasets. Both the query and target sides can involve a combination of text, images, and instructions. Similar to the VQA task, each query has 1,000 candidates, with 1 ground truth and 999 distractors.
Visual Grounding This category includes 1 in-distribution and 3 out-of-distribution datasets, which are adapted from object detection tasks. Queries consist of an instruction, an image, and text referring to a specific region or object within the image. The target may include a cropped image of the object or text describing the same region. Each query includes 1,000 candidates: 1 ground truth and 999 distractors. These distractors may include hard negatives from the same object class, other objects in the image, or random objects from different images.
This task evaluates the model’s ability to recognize and represent the same object or concept across different modalities (image or text) and from varying perspectives. The task involves correctly identifying or referring to specific regions or objects within an image based on the provided instructions or language expressions.
For each evaluation split within the MMEB datasets, 1,000 samples are randomly selected. Further details on dataset processing can be found in Section A.1.
@@ 3.1 Contrastive Training
where “{task_definition}” is a placeholder for a one-sentence description of the embedding task. To enhance the embedding model’s generalizability by better understanding instructions, we have crafted task-specific instructions, as shown in Tables 6, 7, 8 and 9.
@@
where $\mathbb{N}$ denotes the set of all negatives, and $\phi(\mathbf{h}_{q},\mathbf{h}_{t})$ is a function that computes the matching score between query $q$ and target $t$. In this paper, we adopt the temperature-scaled cosine similarity function as follows:
$\phi(\mathbf{h}_{q},\mathbf{h}_{t})=\text{exp}(\frac{1}{\tau}\cos(\mathbf{h}_{q},\mathbf{h}_{t}))$ (3)
$\tau$ is a temperature hyper-parameter.
@@ 4 Experiments
4.1 Experiment Settings
In this paper, we adopt Phi-3.5-V (Abdin et al., 2024) as our backbone VLM, with training conducted via either full model fine-tuning or LoRA. The temperature is set to 0.02. Multiple experimental configurations are explored to assess the impact of key hyperparameters, such as batch size and input length. In the basic setting, we use a batch size of 256, 2,000 training steps, 4 image crops, and a maximum text length of 256 tokens. Detailed ablation studies on these parameters are discussed in Section 4.4.
@@ 4.2 Baselines
CLIP-family: We utilize vision/language encoders such as CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), SigLIP (Zhai et al., 2023), and BLIP2 (Li et al., 2023a) as our baseline. Due to the length limitations of the text encoder, some queries or target text in certain tasks may be truncated. We apply score-level fusion by combining multimodal features using element-wise addition with equal weights ($w_{1}=w_{2}=1$). As found in UniIR (Wei et al., 2023), we do not use instructions, as this was observed to potentially degrade performance.
UniIR: UniIR (Wei et al., 2023) is a unified, instruction-guided multimodal retriever designed to handle eight different retrieval tasks across multiple modalities. The model builds on CLIP and BLIP, employing shallow fusion techniques such as score-level and feature-level fusion to integrate modalities. In this study, we use the BLIP_FF variation as a baseline.
MagicLens: MagicLens (Zhang et al., 2024) is a self-supervised image retrieval model capable of handling open-ended instructions. It utilizes a dual-encoder architecture with shared parameters, initializing the vision and language encoders with either CoCa or CLIP. The model uses a multi-head attention pooler to unify multimodal inputs into a single embedding. For this study, we report results using the CLIP-Large backbone. Since MagicLens requires image inputs, we represent pure text inputs by using the output vector of the text encoder.
E5-V: E5-V (Jiang et al., 2024) is a contemporary model that also leverages vision-language models for multimodal embedding tasks. It proposes a single-modality training approach, where the model is trained exclusively on text pairs. E5-V demonstrated strong performance on three multimodal retrieval tasks. In contrast, we propose MMEB, extending the evaluation to a larger and more diverse set of tasks. Additionally, our model is trained on multimodal pairs, which include various combinations of image and text modalities on both the query and target sides.
@@ 4.3 Main Result
Table 2: Results on the MMEB benchmark. The scores are averaged per meta-task. For detailed scores per dataset, see Table 5.
@@
Baseline Models
@@
LoRA (bs=1024) 54.8 54.9 62.3 79.5 66.5 52.0 60.1
Full Fine-Tuning (bs=1024) 52.8 50.3 57.8 72.3 62.8 47.4 55.9
$\Delta$ Best baseline +7.0 +39.9 +2.2 +17.3 +21.8 +11.6 +17.3
We report Precision@1 for all models in Table 2. It measures the ratio of positive candidates being ranked in the top place for all queries. For Vlm2Vec, both the LoRA and fully fine-tuned variants follow the same setup of 1024 training batch size, 2K training steps and 4 sub-image crops. The LoRA variant uses a rank of 8.
From Table 2, the LoRA version of Vlm2Vec is the best variant, achieving an average precision@1 of 60.1% across all 36 datasets from MMEB. Additionally, it maintains an average precision@1 of 52.0% on 16 out-of-distribution tasks in zero-shot evaluation, suggesting strong generalization ability. This indicates that our model, when well-trained on datasets from diverse task categories, domains, and modality combinations, can effectively follow instructions to align the visual and text spaces and generalize well to unseen tasks. The full fine-tuning variation achieves slightly lower scores than the LoRA version. For a detailed discussion comparing full fine-tuning and LoRA, please refer to Section 4.4.1.
Compared to other baseline models, we observe consistent improvements in our model across all meta-task categories. Notably, our model achieves a 17.3-point improvement (from 42.8 to 60.1) across all 36 MMEB datasets and a 11.6-point increase (from 40.4 to 52.0) on 16 out-of-distribution datasets for zero-shot evaluation. Additionally, unlike the baseline models, which fail to demonstrate reasonable performance across all different task categories, Vlm2Vec achieves relatively strong performance (at least 50%) across all four meta-task categories. This highlights its capability to handle a wide range of multimodal embedding tasks effectively. It’s worth noting that our contemporary work E5-V (Wang et al., 2022a), which is also based on vision-language models, performs much worse than Vlm2Vec. This is due to the fact that E5-V has been trained exclusively on text data without using multimodal data.
@@ 4.4 Result Analysis
To train an effective and generalizable multimodal embedding, various factors need to be considered, ranging from the data to the training setup. In this section, we present detailed ablation studies on these factors. We will discuss two training setups: Full Fine-Tuning vs. LoRA, along with Training parameters, and two topics related to data: Meta-task generalization and Impact of instructions in this section.
@@
Table 3: We compare the performance of fully fine-tuned Vlm2Vec with its LoRA variants at different ranks. LoRA can achieve better performance when the rank is appropriately configured.
@@ 4.4.2 Training parameters
Figure 4: The figures demonstrate the influence of the training setup on Vlm2Vec’s final performance. Here, we examine the effects of training batch size, the number of sub-image crops, and the number of training steps.
@@ 4.4.3 Meta-task generalization
[…]
Figure 5: The figures show the generalization ability of models trained on one meta-task to other unseen meta-tasks. Models trained on retrieval tasks demonstrate better generalization ability because retrieval tasks involve a more diverse combination of text and visual modalities from both the query and target sides.

### v3

@@ Abstract
Embedding models have been crucial in enabling various downstream tasks such as semantic similarity, information retrieval, and clustering. Recently, there has been a surge of interest in developing universal text embedding models that can generalize across tasks (e.g., MTEB). However, progress in learning universal multimodal embedding models has been relatively slow despite its importance and practicality. In this work, we aim to explore the potential of building universal multimodal embeddings capable of handling a wide range of downstream tasks. Our contributions are two fold: (1) we propose MMEB (Massive Multimodal Embedding Benchmark), which covers 4 meta-tasks (i.e. classification, visual question answering, multimodal retrieval, and visual grounding) and 36 datasets, including 20 training datasets and 16 evaluation datasets covering both in-distribution and out-of-distribution tasks, and (2) Vlm2Vec (Vision-Language Model $\rightarrow$ Vector), a contrastive training framework that converts any vision-language model into an embedding model via contrastive training on MMEB. Unlike previous models such as CLIP or BLIP, which encodes text or images independently without any task instruction, Vlm2Vec can process any combination of images and text to generate a fixed-dimensional vector based on the given task instructions. We build a series of Vlm2Vec models on SoTA VLMs like Phi-3.5-V, LLaVA-1.6 and evaluate them on MMEB’s evaluation split. With LoRA tuning, Vlm2Vec can achieve an improvement of 10% to 20% over existing multimodal embedding models on MMEB evaluation sets. We show that VLMs are secretly strong embedding models.
@@ 1 Introduction
- Vlm2Vec: We adopt the pre-trained vision-language models like Phi-3.5-V (Abdin et al., 2024) and LLaVA-1.6 (Li et al., 2024) as the backbone for Vlm2Vec. In contrast to other multimodal embedding models like UniIR (Wei et al., 2023) and MagicLens (Zhang et al., 2024), which rely on late fusion of CLIP (Radford et al., 2021) features, our approach leverages the deep integration of vision and language features within a transformer architecture. There are several advantages to this approach: (1) VLMs are trained on massive multimodal datasets and can handle any combination of images and text, as well as high-resolution images and long text inputs; (2) vision and language features are deeply fused in the transformer model, improving the model’s ability to capture cross-modal relationships; and (3) these models are well-suited for generalizing across diverse tasks, particularly those requiring instruction-following capabilities. These factors make Vlm2Vec an ideal choice for task generalization. We trained Vlm2Vec on the 20 MMEB training datasets using contrastive learning and compared its performance with various baselines.
Following extensive contrastive training, Vlm2Vec can handle any combination of images and text, producing fixed-dimensional vectors. We evaluate Vlm2Vec against a wide array of multimodal embedding models, including CLIP (Radford et al., 2021), BLIP2 (Li et al., 2023a), SigLIP (Zhai et al., 2023), MagicLens (Zhang et al., 2024), UniIR (Wei et al., 2023) and E5-V (Jiang et al., 2024), demonstrating consistent improvements across all task categories. Notably, compared to the best baseline model without fine-tuning, our model achieves a 18.2 point improvement (from 44.7 to 62.9) across all 36 MMEB datasets and a 15.4-point increase (from 41.7 to 57.1) on 16 out-of-distribution datasets for zero-shot evaluation. Compared to the best baseline model with fine-tuning, our model achieves a 15.7 point improvement (from 47.2 to 62.9) across all 36 MMEB datasets and a 14.0-point increase (from 43.1 to 57.1) on 16 out-of-distribution datasets for zero-shot evaluation. Moreover, as a general multimodal representation model, Vlm2Vec can still achieve competitive zero-shot T2I (Text-to-Image) and I2T (Image-to-Text) performance on Flickr30K compared to existing CLIP-like models, as presented in Table 11.
@@ 2.1 Dataset Overview
The embedding models are supposed to compress the query side into a vector and the target candidates into a set of vectors. The candidate with the highest dot-product will be selected as the prediction for evaluation. We measure the Precision@1 to reflect the percentage of top candidate matching the groundtruth. For the number of target candidates, a higher count could increase evaluation costs and hinder rapid model iteration, while a lower count might make the benchmark too simple and prone to saturation. To strike a balance between these extremes, we have chosen 1,000 candidates. Further details about this decision can be found in Section A.2.
MMEB offers a wide range of tasks from various domains, such as common, news, Wikipedia, web, and fashion. The benchmark incorporates diverse combinations of modalities for both queries and targets, including text, images, and text-image pairs. Additionally, tasks are designed to follow different types of instructions. For instance, tasks may involve object recognition (e.g., “Identify the object shown in the image.”), retrieval (e.g., “Find an image that matches the given caption.”), or visual grounding (e.g., “Select the portion of the image that answers the question.”). Examples for each dataset in MMEB are provided in Tables 7, 8, 9 and 10. The diversity in MMEB makes it an ideal testbed for universal embeddings.
@@ 2.2 Meta-task and Dataset Design
Classification The query consists of an instruction, an image, optionally accompanied by related text, while the target is the class label. The number of candidates equals the number of classes.
Visual Question Answering The query consists of an instruction, an image, and a piece of text as the question, while the target is the answer. Each query has 1 ground truth and 999 distractors as candidates.
Information Retrieval Both the query and target sides can involve a combination of text, images, and instructions. Each query has 1 ground truth and 999 distractors as candidates.
Visual Grounding The category is adapted from object detection tasks. The query combines an instruction (e.g., “Select the portion of the image that isolates the object of the given label: red apple”) with the full image. This instruction guides the model to focus on a specific object within the image. Each candidate corresponds to cropped regions (bounding boxes) of the image, including both the object of interest and distractor regions. Each query includes 1,000 candidates: 1 ground truth and 999 distractors. These distractors may include hard negatives from the same object class, other objects in the image, or random objects from different images.
Further details on dataset processing can be found in Section A.1.
@@ 3.1 Contrastive Training
where “{task_definition}” is a placeholder for a one-sentence description of the embedding task. To enhance the embedding model’s generalizability by better understanding instructions, we have crafted task-specific instructions, as shown in Tables 7, 8, 9 and 10.
@@
where $\mathbb{N}$ denotes the set of all negatives, and $\phi(\mathbf{h}_{q},\mathbf{h}_{t})$ is a function that computes the matching score between query $q$ and target $t$. In this paper, we adopt the temperature-scaled cosine similarity function as $\phi(\mathbf{h}_{q},\mathbf{h}_{t})=\text{exp}(\frac{1}{\tau}\cos(\mathbf{h}_{q},\mathbf{h}_{t}))$,where $\tau$ is a temperature hyper-parameter.
@@ 4 Experiments
In this section, we adopt Phi-3.5-V and LLaVA-1.6 as the backbone VLMs, with training conducted via either full model fine-tuning or LoRA. The temperature for the loss function is set to 0.02, with a batch size of 1,024, a maximum text length of 256 tokens, and 2K training steps. The LoRA variant uses a rank of 8. For Vlm2Vec leveraging Phi-3.5-V as the backbone, we configure the number of sub-image crops to 4. For Vlm2Vec using LLaVA-1.6 as the backbone, we resize the input images to a uniform resolution, employing two setups: a high-resolution configuration of 1344 × 1344 and a low-resolution configuration of 336 × 336.
@@
We report Precision@1 for all models in Table 2. It measures the ratio of positive candidates being ranked in the top place for all queries.
@@ 4.1 Baselines
CLIP-family: We utilize vision/language encoders such as CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), SigLIP (Zhai et al., 2023), and BLIP2 (Li et al., 2023a) as our baseline. Due to the length limitations of the text encoder, some queries or target text in certain tasks may be truncated. We apply score-level fusion by combining multimodal features using element-wise addition with equal weights ($w_{1}=w_{2}=1$). We do not use instructions, as they could potentially degrade performance. For more details, please refer to Section 4.3.4.
UniIR: UniIR (Wei et al., 2023) is a unified, instruction-guided multimodal retriever designed to handle eight different retrieval tasks across multiple modalities. The model builds on CLIP and BLIP, employing shallow fusion techniques such as score-level and feature-level fusion to integrate modalities. In this study, we use the CLIP_SF and BLIP_FF variations as baselines.
MagicLens: MagicLens (Zhang et al., 2024) is a self-supervised image retrieval model capable of handling open-ended instructions. It utilizes a dual-encoder architecture with shared parameters, initializing the vision and language encoders with either CoCa or CLIP. The model uses a multi-head attention pooler to unify multimodal inputs into a single embedding. For this study, we report results using the CLIP-Large backbone.
E5-V: E5-V (Jiang et al., 2024) is a contemporary model that also leverages vision-language models for multimodal embedding tasks. It proposes a single-modality training approach, where the model is trained exclusively on text pairs. In contrast, our model is trained on multimodal pairs, which include various combinations of image and text modalities on both the query and target sides.
For all our baselines, we first use their original versions. Additionally, we have fine-tuned both CLIP and OpenCLIP on MMEB training datasets. We adopt the same experimental configurations as Vlm2Vec to ensure a fair comparison. For the remaining baseline models, UniIR and MagicLens also utilize a shallow fusion approach based on CLIP models, with their primary contribution being the datasets they were trained on. E5-V proposes training exclusively on text pairs, making it unsuitable for fine-tuning on our datasets. Therefore, we have not included the fine-tuned versions of these three models in this comparison.
@@ 4.2 Main Result
Table 2: Results on the MMEB benchmark. The scores are averaged per meta-task. For detailed scores per dataset, see Table 6. We include baselines with and without fine-tuning on MMEB training datasets and our models with LLaVA-1.6 and Phi-3.5 backbones. FFT means fully fine-tuned.
@@
Baseline Models (No Fine-tuning on MMEB Training)
@@
Baseline Models (Fine-tuning on MMEB Training)
CLIP-FFT 55.2 19.7 53.2 62.2 47.6 42.8 45.4
OpenCLIP-FFT 56.0 21.9 55.4 64.1 50.5 43.1 47.2
@@
Phi-3.5-V, FFT (bs=1024) 52.8 50.3 57.8 72.3 62.8 47.4 55.9
Phi-3.5-V, LoRA (bs=1024) 54.8 54.9 62.3 79.5 66.5 52.0 60.1
LLaVA-1.6, LoRA (bs=1024,res=336×336) 54.7 50.3 56.2 64.0 61.0 47.5 55.0
LLaVA-1.6, LoRA (bs=1024,res=1344×1344) 61.2 49.9 67.4 86.1 67.5 57.1 62.9
$\Delta$ Best baseline (No Fine-tuning) +16.9 +33.7 +5.6 +20.8 +20.4 +15.4 +18.2
$\Delta$ Best baseline (Fine-tuning) +5.2 +28.0 +12.0 +22.0 +17.0 +14.0 +15.7
From Table 2, the best variant of Vlm2Vec leverages LLaVA-1.6, is trained with LoRA, and processes input images at a relatively high resolution of 1344 × 1344. It achieves an average precision@1 of 62.9% across all 36 datasets from MMEB. Additionally, it maintains an average precision@1 of 57.1% on 16 out-of-distribution tasks in zero-shot evaluation, suggesting strong generalization ability. This indicates that our model, when well-trained on datasets from diverse task categories, domains, and modality combinations, can effectively follow instructions to align the visual and text spaces and generalize well to unseen tasks. It is important to emphasize that LLaVA-1.6 (Li et al., 2024) has a transparent pre-training data recipe and nearly no overlap with our MMEB OOD datasets. This demonstrates that the strong zero-shot results achieved by Vlm2Vec are not attributable to prior exposure of the LLaVA-1.6 backbone to the OOD datasets. When using the same backbone, the full fine-tuning variant achieves slightly lower scores than the LoRA version. For a detailed discussion comparing full fine-tuning and LoRA, please refer to Section 4.3.1.
Compared to other baseline models, with or without fine-tuning on MMEB training data, our model demonstrates consistent improvements. Compared to the best baseline model without fine-tuning, our model achieves a 18.2 point improvement (from 44.7 to 62.9) across all 36 MMEB datasets and a 15.4-point increase (from 41.7 to 57.1) on 16 out-of-distribution datasets for zero-shot evaluation. Compared to the best baseline model with fine-tuning, our model achieves a 15.7 point improvement (from 47.2 to 62.9) across all 36 MMEB datasets and a 14.0-point increase (from 43.1 to 57.1) on 16 out-of-distribution datasets for zero-shot evaluation. Additionally, unlike the baseline models, which fail to demonstrate reasonable performance across all different task categories, Vlm2Vec achieves relatively strong performance (at least 50%) across all four meta-task categories. This highlights its capability to handle a wide range of multimodal embedding tasks effectively.
@@ 4.3 Result Analysis
To train an effective and generalizable multimodal embedding, various factors need to be considered, ranging from the data to the training setup. In this section, we present detailed ablation studies on these factors. We will discuss two training setups: Full Fine-Tuning vs. LoRA, along with Training parameters, and two topics related to data: Meta-task generalization and Impact of instructions.
@@
Table 3: We compare the performance of fully fine-tuned Vlm2Vec with its LoRA variants at different ranks. LoRA can achieve better performance when the rank is appropriately configured. All the models utilize Phi-3.5-V as their backbone.
@@ 4.3.2 Training parameters
Figure 4: The figures demonstrate the influence of the training setup on Vlm2Vec’s final performance. Here, we examine the effects of training batch size, the number of sub-image crops, and the number of training steps. All the models utilize Phi-3.5-V as their backbone.
@@ 4.3.3 Meta-task generalization
We have demonstrated that Vlm2Vec has the potential to transfer to out-of-distribution datasets after being trained on a diverse range of in-distribution datasets, with the instruction-following settings. An interesting question arises as to whether focusing on a specific meta-task can enhance the model’s overall generalizability. We have trained three models, each focused solely on one meta-task (classification, visual question answering, and retrieval). Visual grounding was not included due to the limited number of training datasets. We then evaluated the models’ transferability to other meta-tasks. We refer to these three models as Vlm2Vec ${}_{\text{RET}}$, trained on 8 retrieval tasks, Vlm2Vec ${}_{\text{VQA}}$, trained on 6 visual question answering tasks, and Vlm2Vec ${}_{\text{CLS}}$, trained on 5 classification tasks.
Figure 5 illustrates the generalizability of these three models on unseen meta-tasks. We could observe that Vlm2Vec ${}_{\text{RET}}$ has better generalizablilty on other meta-task, compared with other two models, especially on visual grounding categories. The reason is that retrieval tasks involve a more diverse combination of text and visual modalities from both the query and target sides, which helps the model generalize better to unseen meta-tasks. This observation highlights the benefits of using more diverse tasks in the Vlm2Vec training process.
Figure 5: The figures show the generalization ability of models trained on one meta-task to other unseen meta-tasks. For example, the first subplot compares the performance of Vlm2Vec trained exclusively on VQA datasets with Vlm2Vec trained exclusively on retrieval datasets across the other two meta-task categories: classification and visual grounding. Overall, Vlm2Vec trained on retrieval tasks demonstrate better generalization ability because retrieval tasks involve a more diverse combination of text and visual modalities from both the query and target sides. Vlm2Vec utilizes Phi-3.5-V as its backbone.